### 26 Jan 2024 v2

# The impact of Covid-19 vaccination in Aotearoa New Zealand: a modelling study

Abstract
Aotearoa New Zealand implemented a Covid-19 elimination strategy in 2020 and 2021, which enabled a large majority of the population to be vaccinated before being exposed to the virus. This strategy delivered one of the lowest pandemic mortality rates in the world. However, quantitative estimates of the population-level health benefits of vaccination are lacking. Here, we use a validated mathematical model of Covid-19 in New Zealand to investigate counterfactual scenarios with differing levels of vaccine coverage in different age and ethnicity groups. The model builds on earlier research by adding age- and time-dependent case ascertainment, the effect of antiviral medications, improved hospitalisation rate estimates, and the impact of relaxing control measures. The model was used for scenario analysis and policy advice for the New Zealand Government in 2022 and 2023. We compare the number of Covid-19 hospitalisations, deaths, and years of life lost in each counterfactual scenario to a baseline scenario that is fitted to epidemiological data between January 2022 and June 2023. Our results estimate that vaccines saved 6650 (95% credible interval [4424, 10180]) lives, and prevented 74500 [51000, 115400] years of life lost and 45100 [34400, 55600] hospitalisations during this 18-month period. Making the same comparison before the benefit of antiviral medications is accounted for, the estimated number of lives saved by vaccines increases to 7604 [5080, 11942]. Due to inequities in the vaccine rollout, vaccination rates among Māori were lower than in people of European ethnicity. Our results show that, if vaccination rates had been equitable, an estimated 11–26% of the 292 Māori Covid-19 deaths that were recorded in this time period could have been prevented. We conclude that Covid-19 vaccination greatly reduced health burden in New Zealand and that equity needs to be a key focus of future vaccination programmes.
keywords:
counterfactual modelling, health equity, Māori health, mathematical model, SARS-CoV-2, vaccine
[…]
[inst1]organization=Population Modelling group, National Institute of Water and Atmospheric Research, city=Wellington, country=New Zealand
[…]
[inst2]organization=School of Mathematics and Statistics, University of Canterbury, city=Christchurch, country=New Zealand
[…]
[inst3]organization=Manaaki Whenua, city=Lincoln, country=New Zealand
[…]
[inst4]organization=Department of Engineering Science, University of Auckland, city=Auckland, country=New Zealand
[…]
[inst5]organization=Precision Driven Health, city=Auckland, country=New Zealand
[…]
[inst6]organization=Department of Statistics, University of Auckland, city=Auckland, country=New Zealand
[…]
[inst7]organization=iNZight Analytics Ltd., city=Auckland, country=New Zealand
[…]
author
1 Introduction
Aotearoa New Zealand used a combination of border and community control measures to minimise transmission of SARS-CoV-2 until high vaccine coverage could be achieved baker2020successful ; vattiato2022assessment . Prior to the introduction of the B.1.1.529 (Omicron) variant into the community in January 2022, New Zealand had recorded only around 2.5 confirmed community cases per 1000 people and 0.01 Covid-19 deaths per 1000 people mohgithub . By this time, 77% of the population (90% of those aged over 12 years) had received at least two doses of the Pfizer/BioNTech BNT162b2 vaccine and 27% of the population (35% of adults) had received a third dose. By 1 April 2022, third dose coverage had increased to 51% of the population (66% of adults). Children aged 5 to 11 years became eligible for vaccination on 17 January 2022 and by 1 April, 54% of this age group had received at least one dose and 17% had received two doses mohgithub .
Following transmission from cases in managed isolation facilities into the community in January 2022 douglas2022tracing and subsequent relaxation of border controls, New Zealand experienced a series of Omicron waves. Between 1 January 2022 and 30 June 2023, there were around 465 confirmed cases, 5.5 hospitalisations, and 0.61 Covid-19 deaths per 1000 people mohgithub . Aotearoa New Zealand’s cumulative excess all-cause mortality rate, up to the end of 2022, has been estimated as 0.215 per 1000 people kung2023New , one of the lowest pandemic excess mortality rates in the world.
Achieving high primary series and booster dose coverage before widespread transmission occurred was a key pillar of New Zealand’s pandemic response. Despite good vaccination rates overall, coverage was lower for Māori, partly because of inequities in the vaccine rollout whitehead2021will ; whitehead2022structural ; waitangi2021haumaru . On 1 February 2022, primary series coverage was 76% and 88% for Māori adults under and over 65 years old, respectively. This compares to 88% and 93% for people of European ethnicity (the largest ethnicity group in Aotearoa New Zealand). In addition, Māori are at higher risk of severe illness and death due to Covid-19 steyn2020estimated ; steyn2021maori ; whitehead2023inequities . The age-standardised Covid-19 mortality rate and hospitalisation rate in 2022-23 were approximately 75% higher for Māori than for people of European ethnicity (see Supplementary Material sec. S10).
Globally, it has been estimated that vaccines saved 19–20 million lives in their first year of use watson2022global . However, international estimates cannot simply be mapped onto Aotearoa New Zealand’s population because of its elimination strategy and unique epidemic trajectory. Comparing outcomes with other countries that also pursued an elimination strategy, but experienced waves of transmission with lower vaccine coverage (e.g. Hong Kong smith2022covid ; xie2023resurgence ), suggests that vaccination saved many lives. However, quantitative estimates of the reduction in health burden due to vaccination in New Zealand are currently lacking.
Here, we use a validated mathematical model of SARS-CoV-2 transmission dynamics and health impact in New Zealand to estimate the number of hospitalisations and deaths that were prevented by vaccines between January 2022 and June 2023. This corresponds to the time period in which the Omicron family of variants was dominant esr2023genomics . The model uses robust estimates for the effectiveness of the Pfizer vaccine against infection, severe disease and death caused by Omicron variants khoury2021neutralizing ; cromer2023neutralising ; andrews2022covid . The model builds on our earlier work lustig2023modelling by adding several new features. These include age- and time-dependent case ascertainment, the effect of antiviral medications on fatality rates, improved estimates of the age-dependent infection hospitalisation ratio, the effect of relaxing control measures on contact rates, and the impact of new Omicron sub-variants. The model was used for health system planning, scenario analysis and policy advice for the New Zealand Government in 2022 and 2023.
We fit the model to data on Covid-19 cases, hospitalisations and deaths between January 2022 and June 2023. We then consider counterfactual scenarios in which public health measures, behavioural patterns and resulting time-dependent contact rates were the same as the fitted baseline (factual) scenario, but vaccine coverage was lower. Our results are important for estimating the population-level health benefits of Covid-19 vaccination, supporting evaluation of the cost-effectiveness of Aotearoa New Zealand’s Covid-19 vaccination programme, and for quantifying the impact of inequitable vaccination coverage between population groups. Our findings may be useful in informing preparedness and response to future respiratory pandemic threats.
2 Methods
Mathematical model
We used an age-structured compartment-based model that includes the effects of vaccination, waning immunity and changes in community contact rates over time (see Supplementary Material sec. S1–S4). The population is stratified into 16 five-year age groups. The model includes different levels of immunity to infection and immunity to severe disease and death. Immunity levels depend on the number of vaccine doses received, prior infection status, dominant SARS-CoV-2 variant, and time since last immunising event (see Figure 1). Parameters governing these immunity curves were taken from lustig2023modelling , based on the model of khoury2021neutralizing ; cromer2022neutralising for the relationship between immunity and neutralizing antibody titre.
Figure 1: Average immunity under assumed parameter priors against: (a) infection with BA.2, (b) severe disease or death from BA.2, (c) infection with BA.5 and (d) severe disease or death from BA.5 as a function of time since most recent immunizing event. Graphs show immunity following two doses (blue), three doses (red), 0/1 doses and prior infection with BA.2 (green), two doses and prior infection with BA.2 (magenta) and three doses and prior infection with BA.2 (cyan). Immunity from two or more prior Omicron infections also follows the cyan trajectory, regardless of the vaccination status. Immunity against BA.5 derived from prior infection with BA.5 is assumed to follow the same curves as for immunity against BA.2 derived from prior infection with BA.2. Curves are the median and shaded areas are the 5th and 95th percentiles of 500 random draws from the prior.
The primary series and booster rollout in Aotearoa New Zealand almost exclusively used the Pfizer/BioNTech BNT162b2 vaccine and so we do not attempt to model different effectiveness of different vaccine types. A vaccine pass system was introduced in December 2021 as part of the Covid-19 Protection Framework nzgovt2021covid . Under this system, businesses and venues were required to make proof of vaccination a condition of entry, or to operate with reduced capacity limits. Vaccine mandates were also were introduced for health, disability and education workers and other specified public sector roles nzgovt2021covidb . The vaccine pass system was ended in April 2022 and employment-related vaccine mandates were phased out betwen April and September 2022.
Vaccination rates in the model were set according to Ministry of Health data on the number of 1st, 2nd, 3rd and 4th or subsequent doses given in each age group between 19 February 2021 and 6 June 2023 (see Supplementary Figure S2). At the start of the study period in January 2022, everyone aged over 12 years was eligible for two doses. Eligibility for a third dose was initially limited to over 65-year-olds, but was extended to all adults from 1 February 2022 and to all over-16-year-olds from 7 April 2022. Children aged 5-11 years became eligible for a two-dose paediatric vaccine series from 17 January 2022. In June 2022, over-50-year-olds became eligible for a fourth dose. From April 2023, all over-30-year olds became eligible for an additional dose 6 months after their more recent dose. In all cases, people outside the age criteria are eligible if they have a specified health condition. In April 2023, the Pfizer bivalent BA.4/5 vaccine replaced the original formulation for use as a booster dose. However, since this was designed primarily to counter the effects of antigenic evolution, we assume this had the same effectiveness parameters as the original vaccine did against earlier Omicron strains.
The model has previously been fitted to Aotearoa New Zealand data on Covid-19 cases, hospital admissions and deaths, and used to model the impact of the BA.5 subvariant that caused a wave in July 2022 lustig2023modelling . The model used in this study additionally included age- and time-dependent case ascertainment and an adjusted age-dependent infection-hospitalisation ratio to better match the observed age structure in hospital admissions (see Supplementary Material sec. S8). We also made additions to the model to account for changes in the public health response and the impact of new Omicron sub-variants described below.
On 13 September 2022, the New Zealand Government ended the Covid-19 Protection Framework nzgovt2021covid , which meant that mask requirements in public settings such as retail and public transport were lifted, and isolation requirements for household contacts were ended. To model the effect of this policy change, we included an increase in contact rates in the model, with the magnitude of the effect fitted to subsequent data (see Supplementary Material sec. S6). This is in addition to the increase in contact rates and within-age-group mixing in March-April 2022, representing relaxation of non-pharmaceutical interventions and associated precautionary behaviour, as previously modelled lustig2023modelling .
Around the same time the Covid-19 Protection Framework ended, the eligibility criteria for antiviral treatments for Covid-19 were broadened to include everyone aged over 65 years, Māori and Pacific people aged over 50 years, and people with specified comorbidities. This was followed by a significant increase in the number of prescriptions for antivirals (Supplementary Figure S5). To model this, we assumed that the infection-fatality ratio at time $t$ was a linearly decreasing function of the proportion of notified cases at time $t$ who filled a prescription for either Paxlovid or molnupiravir, the two Covid-19 antiviral treatments in widespread use in New Zealand during this time period (see Supplementary Material sec. S8). The effect size was fitted to subsequent data. We did not attempt to distinguish between different effect sizes for different antivirals as the data were insufficient for this. We did not model any effect of antivirals on the infection-hospitalisation rate, partly because in some cases antiviral prescriptions were only filled on or after hospital admission.
In November 2022, a mixture of Omicron lineages (most notably CH.1.1 and BQ.1.1) sharing similar sets of genetic mutations displaced BA.5 as the most common variant esr2023genomics . To model the growth advantage of these subvariants, we assumed that, on 15 November 2022, a new combination of functionally equivalent immune evasive subvariants became dominant. This is a highly simplified model: we did not attempt to model these subvariants individually, but instead assumed that their net effect can be captured via a reduction in the level of population immunity. We implemented this via the same model mechanism as for the BA.5 subvariant that displaced BA.2 as the dominant variant in July 2022 lustig2023modelling , with an effect size parameter that was manually calibrated using subsequent data (see Supplementary Material, sec. S7).
We fitted the model to data on confirmed Covid-19 cases, hospitalisations and deaths between 1 February 2022 and 13 August 2023. Hospitalisations were defined to be those categorised by the Ministry of Health as receiving hospital treatment for Covid-19. Deaths were defined to be those where the cause of death was classified as “COVID underlying” or “COVID contributory”. The Ministry of Health also publishes the number of deaths that occurred within 28 days of a positive Covid-19 test. The definition that we use is a more accurate estimate of the impact of Covid-19 as it excludes incidental deaths (i.e. those where the cause of death was found to be not Covid-19-related). Fitting was done with an approximate Bayesian computation (ABC) rejection algorithm on a subset of model parameters including time- and age-dependent contact rates, testing and clinical severity parameters, waning rate, and antiviral effect size. This results in a set of 150 model realisations, with each one representing a combination of parameter values sampled from the approximate joint posterior distribution. For all results, we report 95% credible intervals (CrI) corresponding to the 2.5th and 97.5th percentiles of the sample (see Supplementary Material sec. S9 for details).
For the main analysis, we used the Ministry of Health’s Health Service User (HSU) dataset (see Supplementary Table S2) for population size by age and prioritised ethnicity. However, this dataset underestimates the size of the Māori population and has a different age profile than official population counts and estimates statsnz_hsu . We therefore also ran a sensitivity analysis using StatsNZ’s annual population estimates tepou2021dhb , in which the Māori population size is 11% larger than in the HSU data (see Supplementary Material sec. S5 for details).
Counterfactual scenarios
Figure 2: Cumulative number of vaccine doses per capita in the baseline scenario (representing actual data up to 6 June 2023, blue) and counterfactual scenarios: vaccination rates set to a proportion $p_{v}=0.9$ of actual vaccination rates at all ages (red); vaccination rates at all ages set to actual vaccination rates in the 20-25-year-old group (yellow); vaccination rates set to actual vaccination rates for the European/other (purple) and Māori (green) ethnicities. The scenarios with no vaccination and no vaccination in under-60-year-olds are not shown as they are either zero or equal to the baseline (blue) curves. Each panel shows a different age group. Graphs only show total number of doses for ease of visualisation but in the model these are broken down into 1st, 2nd, 3rd, and 4th or subsequent doses.
We ran model simulations with the same set of parameter combinations produced by the baseline fitting procedure, but with the number of daily $d^{\mathrm{th}}$ vaccine doses $v_{id}(t)$ in age group $i$ at time $t$ modified to represent a given counterfactual scenario. For each accepted parameter combination, we considered the following counterfactual scenarios: (1) no vaccination; (2) vaccination rates set to a proportion $p_{v}$ of actual vaccination rates at all ages; (3) no vaccination of under-60-year-olds; (4) vaccination rates for all adults set to actual vaccination rates in the 20-25-year-old group (resulting in a reduction in coverage particularly for older adults); (5-6) vaccination rates set to actual vaccination rates for European/other and Māori ethnicities respectively at all ages (see Figure 2). Where vaccination rates were changed by a multiplicative factor, or set equal to those of another age or ethnicity group, this change was applied to the time series of daily 1st, 2nd, 3rd, and 4th or subsequent doses (see Supplementary Table S3 for details). We also ran the baseline scenario and scenario (1) but with no antiviral medications in the model. This was to enable a secondary comparison of outcomes with and without vaccination before the impact of antivirals on the death rate is accounted for.
In each scenario, we assumed that the time-dependent contact rates, mixing between age groups, timing and characteristics of new variants were the same as in the fitted factual scenario. In each scenario, we calculated total number of infections, hospital admissions, deaths and years of life lost (YLL) due to Covid-19, and the peak hospital occupancy.
Population-level YLL were calculated using New Zealand cohort life tables published by StatsNZ statsnz_cohort_life_tables . The life tables provide life expectancy estimates by age, sex and year of birth. We used the model output for the total number of Covid-19 deaths in each five-year age group and imputed the number of model deaths in one-year age groups and by sex according to the distribution of actual Covid-19 deaths. We then multiplied this by the life expectancy estimate for people of that age and sex in 2022, and calculated total YLL by summing across all age groups. We only calculated YLL at the population level and did not attempt to calculate YLL for specific ethnicity groups.
To estimate the number of Māori hospitalisations and deaths that could have been prevented if Māori vaccination rates had been equal to those of European/other ethnicity, we first calculated the model death (or hospitalisation) rates in age group $i$ in the scenarios with European/other vaccination rates ($E_{i}$) and Māori vaccination rates ($M_{i}$) aggregated over the modelled time period. We then applied the relative difference between these rates to the actual number of Māori deaths (or hospitalisations) that occurred in each age group in the same time period, and summed over all age groups:
[…]
The time period considered for counterfactual scenarios was 1 January 2022 to 30 June 2023, which is the first 18 months of the Omicron period. We did not consider hospitalisations and deaths prior to 1 January 2022 and did not attempt to model the dynamics of the Delta outbreak that occurred in Auckland between August 2022 and January 2023 plank2022using . The national vaccine rollout was still underway during the period and intensive non-pharmaceutical interventions were used to suppress transmission and contain the outbreak to the Auckland metropolitan area up to December 2022. Over 99% of all cases and around 98% of all deaths reported up to 30 June 2023 occurred after 1 January 2022.
Data and fully documented code to reproduce the results in this article is publicly available at https://github.com/SamikDatta/covid19_vaccination/tree/main.
3 Results
Baseline model fit
The model provided a reasonable fit to the time series of cases, hospitalisations and deaths between January 2022 and June 2023 at the aggregate level (Figure 3) and within ten-year age bands (Supplementary Figures S8–S9). This time period included three distinct waves dominated by different Omicron subvariants: the BA.1/BA.2 wave that peaked in March 2022, the BA.5 wave that peaked in July 2022, and the predominantly CH.1.1/BQ.1.1 wave that peaked in December 2022. There was also a smaller wave that peaked in April 2023 and was dominated by a mixture of recombinant XBB lineages esr2023genomics . However, unlike the two preceding waves, this wave did not require specific model assumptions about a more transmissible or immune evasive variant, but was largely captured by the default model assumptions about continuous background waning of immunity.
The fitted baseline model had 3163 (95% CrI [2169, 4561]) deaths and 28800 [22200, 35600] hospital admissions between 1 January 2022 and 30 June 2023. The actual numbers for this period were 3196 deaths and 28763 admissions. Our model estimated that Covid-19 was responsible for 39900 [27600, 57300] YLL in this time period, an average of approximately 13 YLL per Covid-19 fatality.
Figure 3: Results for the baseline (factual) scenario showing: (a) new daily infections per 10,000 people; (b) new daily cases; (c) new daily hospital admissions; (d) hospital occupancy; (e) daily deaths. Graphs show the curvewise 95% CrI (grey shaded area) and a random sample of model trajectories (solid gray lines). Model was fitted to data (blue points) up to 13 August 2023; solid blue curves show the moving average of the data in a 7-day window for cases and hospital occupancy, a 14-day window for admissions, and a 21-day window for deaths.
Counterfactual scenarios
In the counterfactual scenario with no vaccination, there were a total of 9953 [6616, 14824] deaths, 115100 [79500, 174800] YLL, and 74300 [56500, 90900] hospital admissions between 1 January 2022 and 30 June 2023 (see Table 1). This represents an additional 6650 [4424, 10180] deaths, 74500 [51000, 115400] YLL, and 45100 [34400, 55600] admissions relative to the baseline scenario (see Supplementary Table S5). An alternative basis for estimating the effect of vaccination is to compare model outcomes with and without vaccination in a scenario with no antivirals. Under this comparison, the scenario with no vaccination had an additional 7604 [5080, 11942] deaths and 82400 [56100, 129300] YLL (see Supplementary Table S5).
These additional health impacts were concentrated in the first Omicron wave in February-March 2022 (see Figure 4), which infected around 50% more people than in the baseline scenario and had higher average severity since all infections were in immune naive individuals. Peak hospital occupancy in this wave with no vaccines was 5542 [3695, 7526]. This is around 7 times higher than in the baseline scenario, a level that would have completely overwhelmed hospital capacity (see Discussion). Additional health impacts continued to accumulate over the remainder of the study period, albeit at a slower rate (Figure 4).
In the no vaccination scenario, the attack rate (proportion of the population with at least one infection) in the over-75-year-old group was 52% [43%, 62%]. This was significantly lower than attack rates in younger groups, which were over 90% for under-60-year-olds. This means that, in the no vaccination scenario, there was still as a significant immune naive population in elderly groups at the end of the simulation period. As a consequence, the health burden prevented by vaccination will continue to increase over time.
Scenarios (2)–(4) had varying levels of impact intermediate between the baseline and no vaccination scenarios (Table 1). Of these, scenario (3) with no vaccination in under-60-year-olds had the biggest impacts, most notably on hospital admissions and YLL. This is due to shallower age gradient in hospitalisation risk relative to fatality risk (see Supplementary Table S2) and an increase in fatalities under-60-year-olds where life expectancy is higher.
Scenario
Infections       (millions)
Admissions       (thousands)
Deaths
YLL              (thousands)
Peak occupancy
Actual
$28.8$
$3196$
$33.9^{*}$
$1016^{**}$
Model scenarios
(0) Baseline
5.83 [5.12, 6.77]
28.8 [22.2, 35.6]
3163 [2169, 4561]
39.9 [27.6, 57.3]
833 [549, 1266]
(0a) No AVs
5.83 [5.12, 6.77]
28.8 [22.2, 35.6]
3818 [2576, 5427]
45.3 [31.0, 64.8]
833 [549, 1266]
(1) No vaccine
7.27 [6.52, 8.29]
74.3 [56.5, 90.9]
9953 [6616, 14824]
115.1 [79.5, 174.8]
5542 [3965, 7256]
(1a) No vaccine or AVs
7.27 [6.52, 8.29]
74.3 [56.5, 90.9]
11457 [7671, 17341]
128.4 [88.4, 194.0]
5542 [3965, 7256]
(2) 10% drop in rates
5.96 [5.25, 6.91]
33.1 [25.6, 40.7]
3788 [2569, 5519]
46.8 [32.2, 68.2]
1036 [693, 1388]
(3) No vaccine in U60s
7.02 [6.27, 8.01]
43.8 [33.2, 52.5]
3937 [2706, 5695]
59.5 [41.0, 87.6]
2845 [2076, 3620]
(4) 20-25-year-old rates
6.03 [5.29, 7.00]
33.2 [25.6, 41.0]
3906 [2660, 5666]
47.2 [32.5, 68.0]
916 [593, 1313]
(5) Euro/other rates
5.80 [5.09, 6.74]
28.4 [21.8, 35.1]
3100 [2127, 4460]
39.2 [27.2, 56.4]
817 [531, 1260]
(6) Māori rates
6.06 [5.33, 7.02]
31.9 [24.7, 39.4]
3522 [2410, 5107]
43.7 [30.3, 63.4]
1011 [681, 1312]
Table 1: Model results (median and 95% CrI) in each scenario for the total number of infections, hospital admissions, deaths, and years of life lost (YLL), and the peak hospital occupancy, between 1 January 2022 and 30 June 2023. Scenarios are: (0) baseline (actual vaccination rates); (0a) no antivirals (actual vaccination rates); (1) no vaccination; (1a) no vaccination or antivirals; (2) vaccination rates set to a proportion $p_{v}=0.9$ of actual vaccination rates at all ages; (3) no vaccination of under-60-year-olds; (4) vaccination rates at all ages set to actual vaccination rates in the 20-25-year-old group; (5-6) vaccination rates set to actual vaccination rates for European/other and Māori ethnicities respectively at all ages. $*{}^{*}$stimated from actual deaths using cohort life tables via the same method as for model YLL calculations. ${}^{**}$ncludes some incidental hospitalisations (i.e. patients who were positive for Covid-19 but not receiving treatment for Covid-19).
Scenario (2), which was a 10% reduction in vaccination rates at all ages, and scenario (4), in which vaccination rates at all ages were set to those of 20-25-year-olds had very similar outcomes. The reduction in vaccination in older groups in scenario (4) was partially offset by the fact that 20-25-year-olds generally received their primary series and third dose somewhat later than older groups (see Figure 2). The timing of the first wave in February-March 2022 therefore meant that older groups were less affected by waning immunity in scenario (4) than in the baseline scenario. With different timing, the additional health impact in this scenario could have been even larger, emphasising the importance of achieving high vaccine coverage in older age groups.
Relative differences among scenarios in the number of infections were smaller ($<$25% above baseline in all scenarios). This shows that, over the 18-month time period considered, the primary benefit of vaccination in the model was to reduce the risk of severe disease and death rather than to prevent transmission. The relative differences in hospital admissions and deaths of all scenarios compared to baseline are shown in Figure 4.
Figure 4: Cumulative number of infections, hospital admissions and deaths over time in the baseline scenario (red) and the scenarios with no vaccination (green), no vaccination or antivirals (purple), and no vaccination of under 60-year-olds (yellow). Graphs show the median (solid curves) and 95% CrI (shaded areas) for each scenario. Results for the other scenarios considered are not shown here as they are relatively close the baseline scenario (see Supplementary Figure S10). Note that, for infections and admissions, the results for the ‘No vaccine’ and ‘No vaccine or antivirals’ scenarios are identical as the model assumes that antivirals only affect the fatality rate.
Scenario (5) in which European/other vaccination rates were applied to the whole population had 64 [43, 91] fewer deaths and 500 [400, 600] fewer hospitalisations than baseline (Figure 4 and Supplementary Table S5). Scenario (6) in which Māori vaccination rates were applied to the whole population had a death rate that was 8.0 [5.4 12.2] per 100,000 higher and a hospitalisation rate 68.7 [52.1, 82.7] per 100,000 higher than scenario (5). These differences account for 13–29% and 14–23% of the observed differences in age-standardised Covid-19 mortality and hospitalisation rates respectively (see Supplementary Table S4).
Applying the relative differences in age-specific death and hospitalisation rates between scenarios (5) and (6) to the actual number of Māori deaths and hospitalisations suggests that if Māori vaccine uptake had been at the same level as European/other, 32 [30, 36] deaths and 457 [412, 505] hospitalisations could have been prevented.
In the sensitivity analysis using StatsNZ projections instead of the HSU dataset for population denominators, the results for scenarios (0)–(4) (Supplementary Table S6) were similar to the main analysis shown in Table 1. However, scenario (6) in which Māori vaccination rates were applied had significantly worse outcomes than in the main analysis. The estimated numbers of preventable Māori deaths and hospitalisations in the sensitivity analysis were 77 [72, 84] and 998 [914, 1104] respectively.
Considering both the main and the sensitivity analysis, our results suggest that 11–26% of the 292 Māori Covid-19 deaths and 10–22% of the 4616 Māori hospitalisations that were recorded during the time period could have been prevented had vaccination rates been equitable.
4 Discussion
We have used a validated mechanistic model of Covid-19 transmission dynamics and health impact in Aotearoa New Zealand to explore outcomes in a range of counterfactual scenarios where public health measures and time-varying contact rates were the same as the baseline (factual) scenario, but vaccine coverage was lower. Our results showed that vaccines saved an estimated 4600–9500 lives and prevented 34000–58000 hospitalisations between January 2022 and June 2023, the first 18 months of widespread community transmission. These estimates will continue to increase over time as the model estimated that a significant proportion of the elderly population remained uninfected in June 2023.
Benefits cannot always be attributed exclusively to a single intervention such as vaccines, particularly when interventions act multiplicatively to reduce risk. In our model, this applies to the benefits of antiviral medications and vaccines. To address this, we compared scenarios with and without vaccination and antivirals. When the comparison of outcomes with and without vaccination was made before the benefit of antivirals on the death rate is taken into account, the estimated number of lives saved by vaccines increased to 5300–10800.
These estimates only include direct health impacts of Covid-19 and do not account for indirect effects, such as a potential increase in the death rate from Covid-19 or other causes due to an overloaded healthcare system. Similarly, the modelled hospital occupancy does not take account of any change in admission thresholds that may have occurred due to capacity constraints. It should therefore be interpreted as an estimate of the demand for hospital treatment rather than a prediction of realised occupancy.
Vaccination may have prevented other adverse outcomes, such as symptomatic disease leading to attendance at primary care, lost productivity, and incidence of Long Covid notarte2022impact ; tran2023efficacy . However, we only focused on hospitalisations and deaths and did not consider other outcomes in the model due to a lack of relevant data.
Adverse events following immunisation with COVID-19 vaccines in Aotearoa New Zealand are monitored by the statutory regulator Medsafe. Detailed reports are available at medsafe_adverse . Serious adverse events are extremely rare; however we have not attempted a formal risk-benefit analysis here.
A global modelling study estimated that vaccines prevented approximately 20 million deaths up to December 2021 watson2022global , which equates to around $2.5$ deaths prevented per 1000 primary series doses. Another analysis estimated that vaccines prevented 105,900 deaths in England up to August 2021 ukhsa2021vaccine , equivalent to $1.4$ per 1000 primary series doses. By comparison our results equate to between $0.54$ and $1.28$ deaths prevented per 1000 primary series doses (of which there were $8.45$ million during the simulation period).
There are several reasons why the number of deaths prevented per dose would be lower in the context of our study. In most countries, the initial vaccine rollout prioritised high-risk groups and took place during a period of high global community transmission of the Alpha, Beta and Delta variants. It is to be expected that there are diminishing returns from expanding the rollout to lower-risk groups, even though these are almost certainly still net beneficial. The Omicron variant is less pathogenic than earlier variants of concern nyberg2022comparative and effective treatments have improved over time. Finally, although vaccine effectiveness against severe disease caused by Omicron has remained high, effectiveness against transmission is significantly reduced. Globally, as well as in Aotearoa New Zealand, cumulative attack rates were far higher by June 2023 than in December 2021.
Counterfactual modelling is a valuable tool to understand and learn from the potential consequences of alternative scenarios with the benefit of hindsight provided by subsequent data mishra2021comparing ; binny2021early ; imai2023quantifying . However, it has important limitations.
In reality, in a scenario with low vaccine coverage, the occurrence of higher hospitalisation and death rates would have prompted different policy actions and behavioural responses gimma2022changes , which would have altered the transmission dynamics compared to the counterfactual scenarios presented here. Therefore our results should not be interpreted as predictions of what would have happened if Aotearoa New Zealand had not achieved as high a vaccine coverage as it did. Nevertheless, in the absence of vaccination, non-pharmaceutical interventions would only have delayed rather than prevented infections. It is likely that the large majority of unvaccinated individuals would eventually have been infected, meaning a comparable number of hospitalisations and deaths as in the modelled counterfactual would have accrued, albeit over a longer time period. Our results are therefore a valid estimate of the reduction in health burden that is directly attributable to vaccines. We also note that the need for a more stringent policy or behavioural response would have incurred additional direct and indirect costs.
The model for transmission dynamics does not include ethnicity as a variable. In estimating the number of Māori deaths that could have been prevented if vaccination rates had been equitable, we assumed that the per capita death rate for a subgroup with a given vaccination rate would be equal to the per capita death rate if that vaccination rate was applied to the whole population. This is a simplistic assumption that ignores the differential effect of vaccination on transmission between and within ethnicities. Our estimates apply only to the period from 1 January 2022 to 30 June 2023. They do not include vaccine-preventable health impacts that disproportionately affected Māori during the 2021 Delta outbreak tewhatuora_dashboard ; smith2022covid ; summers2023improvements .
We found that estimates for the number of preventable Māori hospitalisations and deaths were larger when using StatsNZ population estimates than when using the HSU data for population denominators. It is recognised that the HSU data underestimates total Māori population size and has biases in the age and sex structure as it only includes those with recent engagement with health services statsnz_hsu . Therefore the results that use the StatsNZ population estimates may be more accurate for Māori impacts. The significant gap between these point to the need for consistent collection of high-quality ethnicity data, accurate population size estimates, and explicit consideration of the uncertainty due to the differences in key data sources.
Our results show that inequitable vaccine coverage is one contributing factor to the higher Covid-19 hospitalisation and mortality rates experienced by Māori. Other factors likely include increased exposure to infection due to high rates of employment in high-contact occupations, larger average household size and lower-quality housing mcleod2020covid ; harvey2023summary , high rates of comorbidities steyn2020estimated ; yon2014cohort ; gurney2020inequity , poor access to healthcare whitehead2022structural , and a health system that has historically underserved Māori waitangi2021haumaru . All these factors are worthy of attention and further study. Including the effects of ethnicity in dynamic epidemiological models is challenging due to the paucity of data and need for modelling assumptions about contact rates between and within ethnic groups. This a key area needing further research in Aotearoa New Zealand.
Our results cannot be used to estimate outcomes if there had been widespread transmission of SARS-CoV-2 in Aotearoa New Zealand earlier in the pandemic when fewer effective treatments were available. Some pre-Omicron variants were more pathogenic nyberg2022comparative ; davies2021increased and it is likely that health impacts would have been larger if uncontrolled transmission of these variants had occurred in an unvaccinated population. We did not attempt to model counterfactual scenarios based on alternative vaccination policies, such as absence of the vaccine pass system and employment-related vaccine mandates. This is because there is no quantitative or qualitative data to inform modelling assumptions about the effect these policies would have had on vaccine update or on contact rates of unvaccinated individuals.
Our analysis is retrospective and cannot be used to predict the health benefits of future booster rollouts or changes to vaccine eligibility. These will depend on the future prevalence and clinical severity of SARS-CoV-2 variants, the marginal increase in protection offered by additional doses relative to existing levels of hybrid immunity, and how well matched the vaccines are to dominant circulating strains fabiani2023relative ; kirsebom2023effectiveness .
The uncertainty estimates we have provided arise from the range of parameter values for which model outputs are approximately consistent with epidemiological data in the baseline scenario. There may be additional uncertainty arising from potential model misspecification, omission of relevant mechanisms, or population heterogeneity, which were not accounted for.
Aotearoa New Zealand’s Covid-19 pandemic response used a combination of border controls, physical distancing measures including periods of stay-at-home orders, financial support for those unable to work, testing and contact tracing, mass masking, vaccination, and antiviral treatments baker2020successful ; cumming2022going ; binny2022real . This response delivered one of the lowest pandemic mortality rates of any country in the world cmi2023mortality . Given that non-pharmaceutical interventions could only delay rather than prevent widespread transmission, it is clear that vaccination was an essential component of this response that saved thousands of lives. Immunity provided by existing vaccine coverage and future vaccination will continue to be an important factor in reducing the Covid-19 health burden. This means the number of lives saved by Covid-19 vaccines in Aotearoa New Zealand will further increase over time. Ensuring equitable access to vaccination and vaccine coverage should be a priority for future public health or pandemic responses.
Acknowledgements
The authors acknowledge the role of the New Zealand Ministry of Health, StatsNZ, and the Institute of Environmental Science and Research in supplying data in support of this work. The authors are grateful to Fiona Callaghan, Nigel French, Emily Harvey, Markus Luczak-Roesch, Dion O’Neale, Matt Parry, Patricia Priest, the Covid-19 Modelling Government Steering Group and two anonymous reviewers for feedback on earlier versions of this manuscript. The authors acknowledge the contributions of Rachelle Binny, Shaun Hendy, Kannan Ridings, Nicholas Steyn and Leighton Watson to previous models from which this model was developed. This work was partly funded by the New Zealand Department of the Prime Minister and Cabinet and Ministry of Health.